# Covariant Quantum Error-Correcting Codes with Metrological Entanglement Advantage

### v1

@@ Acknowledgements.
where $\lambda_{ij}$ are the components of a density operator, and $D_{P}(\Lambda+\mathcal{B},\Lambda)\leq\epsilon$, where $\Lambda(\rho)=\sum_{i,j}\lambda_{ij}\text{Tr}(\rho)|i\rangle\langle j|$ and $(\Lambda+\mathcal{B})(\rho)=\Lambda(\rho)+\sum_{i,j}\text{Tr}(\rho B_{ij})|i\rangle\langle j|$.
This theorem in fact has an interpretation from the complementary channel point of view, as described in Ref. [18]. More specifically, consider the noise channel $\mathcal{N}(\rho)=\sum_{i}K_{i}(\rho)K_{i}^{\dagger}$, whose complementary channel is $\hat{\mathcal{N}}(\rho)=\sum_{i,j}\text{Tr}(K_{j}\rho K_{i}^{\dagger})|i\rangle\langle j|=(\Lambda+\mathcal{B})(\rho)$, where $|i\rangle\langle j|$ acts on the environment Hilbert space. If we have perfect Knill-Laflamme conditions, the complementary channel of the noise becomes $\hat{\mathcal{N}}(\rho)=\sum_{i,j}\lambda_{ij}|i\rangle\langle j|:=\lambda_{0}=\Lambda(\rho)$, which is a fixed density matrix $\lambda_{0}$ independent of the input $\rho$. In other words, the condition for having a perfect quantum error-correcting code is equivalent to the condition that the environment gains no information from the code space. And we indeed see that the AQECC generalization of the Knill-Laflamme conditions is a perturbation from the complementary channel point of view, where the environment can gain some $\epsilon$-small information from the code space. Note that, in the original theorem in Ref. [18], Bures distance $D_{B}(\mathcal{N},\mathcal{M})$ is used as the measure of channel distance, while here we opt for using the purified distance $D_{P}(\mathcal{N},\mathcal{M})$, where $D_{B}(\mathcal{N},\mathcal{M})^{2}\leq D_{P}(\mathcal{N},\mathcal{M})$ can be obtained using Eq. (S2).
Here we consider the noise channel $\mathcal{N}(\rho)=\sum_{i=1}^{n_{d}}K_{i}(\rho)K_{i}^{\dagger}$, where $n_{d}=(2s+1)^{2d}$ is the number of independent Kraus operators. We assume $K_{j}$ is at most $d$-local, and that all $K_{j}$ act on the same set of $d$ qudits, so $K_{i}^{\dagger}K_{j}$ is at most $d$-local as well. We are also interested in the regime $J\sim N^{a}$ and $(M_{\mathrm{max}}-M_{\mathrm{min}})\sim N^{b}$. From Lemma 1, for $|J,n\rangle,|J,m\rangle\in\mathfrak{C}$ and $n\neq m$, we have $\langle J\!,\!{n}|K_{i}^{\dagger}K_{j}|J\!,\!{m}\rangle=0$ if $\Delta\geq 2sd+1$. Now assuming we take $\lambda_{ij}=\langle J\!,\!M_{\mathrm{min}}|K^{\dagger}_{i}K_{j}|J\!,\!M_{\mathrm{min}}\rangle$, we then have $\Lambda(\rho)=\sum_{i,j=1}^{N_{d}}\lambda_{ij}\text{Tr}[\rho]|i\rangle\langle j|$ and $(\Lambda+\mathcal{B})(\rho)=\sum_{i,j=1}^{N_{d}}\text{Tr}[K_{j}\rho K_{i}^{\dagger}]|i\rangle\langle j|$, where the inputs of both channels are from the logical space.
@@
For any state $|\psi\rangle=\sum_{x,y}A_{xy}|x\rangle\otimes|y\rangle$, if we consider the operators $A=\sum_{x,y}A_{xy}|x\rangle\langle y|$ and $B:=A^{T}=\sum_{x,y}A_{xy}|y\rangle\langle x|$ and the state $|\Phi\rangle=\sum_{x=1}^{2^{k}}|x\rangle\otimes|x\rangle$, then we can express $|\psi\rangle=(A\otimes\mathrm{id})|\Phi\rangle=(\mathrm{id}\otimes B)|\Phi\rangle$. We can therefore rewrite $\mathcal{E}\otimes\mathrm{id}(|\psi\rangle\langle\psi|)=(\mathrm{id}\otimes B)(\mathcal{E}\otimes\mathrm{id})(|\Phi\rangle\langle\Phi|)(\mathrm{id}\otimes B^{\dagger})$ for $\mathcal{E}=\Lambda$ and $\mathcal{E}=\Lambda+\mathcal{B}$. Moreover,
$(\Lambda+\mathcal{B})\otimes\mathrm{id}(|\Phi\rangle\langle\Phi|)=\sum_{x=1}^{2^{k}}\sum_{i,j=1}^{n_{d}}\langle J\!,\!M_{x}|K_{i}^{\dagger}K_{j}|J\!,\!M_{x}\rangle\cdot|i\rangle\langle j|\otimes|x\rangle\langle x|:=\sum_{x=1}^{2^{k}}\sum_{i,j=1}^{n_{d}}\Sigma^{x}_{ij}\cdot|i\rangle\langle j|\otimes|x\rangle\langle x|,$ (S14)
$\Lambda\otimes\mathrm{id}(|\Phi\rangle\langle\Phi|)=\sum_{x=1}^{2^{k}}\sum_{i,j=1}^{n_{d}}\langle J\!,\!M_{\mathrm{min}}|K_{i}^{\dagger}K_{j}|J\!,\!M_{\mathrm{min}}\rangle\cdot|i\rangle\langle j|\otimes|x\rangle\langle x|=\sum_{x=1}^{2^{k}}\sum_{i,j=1}^{n_{d}}\lambda_{ij}\cdot|i\rangle\langle j|\otimes|x\rangle\langle x|~{},$ (S15)
where we have used Lemma 1 and denoted $\Sigma^{x}_{ij}=\langle J\!,\!M_{x}|K_{i}^{\dagger}K_{j}|J\!,\!M_{x}\rangle$.
Note that the normalization condition $\langle\psi|\psi\rangle=\sum_{x}\langle x|B^{\dagger}B|x\rangle=1$, so we can interpret $\langle x|B^{\dagger}B|x\rangle$ as a probability. The fidelity can be rewritten as
@@
$=f\left(\sum_{x=1}^{2^{k}}\langle x|B^{\dagger}B|x\rangle\cdot\sum_{i,j=1}^{n_{d}}\Sigma_{ij}^{x}|i\rangle\langle j|\otimes\frac{B|x\rangle\langle x|B^{\dagger}}{\langle x|B^{\dagger}B|x\rangle},\sum_{x=1}^{2^{k}}\langle x|B^{\dagger}B|x\rangle\cdot\sum_{i,j=1}^{n_{d}}\lambda_{ij}|i\rangle\langle j|\otimes\frac{B|x\rangle\langle x|B^{\dagger}}{\langle x|B^{\dagger}B|x\rangle}\right)$
$\geq\sum_{x=1}^{2^{k}}\langle x|B^{\dagger}B|x\rangle\cdot f\left(\sum_{i,j=1}^{n_{d}}\Sigma_{ij}^{x}|i\rangle\langle j|\otimes\frac{B|x\rangle\langle x|B^{\dagger}}{\langle x|B^{\dagger}B|x\rangle},\sum_{i,j=1}^{n_{d}}\lambda_{ij}|i\rangle\langle j|\otimes\frac{B|x\rangle\langle x|B^{\dagger}}{\langle x|B^{\dagger}B|x\rangle}\right)$
$=\sum_{x=1}^{2^{k}}\langle x|B^{\dagger}B|x\rangle\cdot f\left(\sum_{i,j=1}^{n_{d}}\Sigma_{ij}^{x}|i\rangle\langle j|,\sum_{i,j=1}^{n_{d}}\lambda_{ij}|i\rangle\langle j|\right)~{},$ (S16)
@@
We will use the Fuchs-van de Graaf inequality Eq. (S1) to bound the fidelity via the trace norm. To this end, we would like to bound $\|\Sigma_{x}-\lambda_{0}\|_{1}$, where $\Sigma_{x}:=\sum_{i,j=1}^{N_{d}}\Sigma_{ij}^{x}|i\rangle\langle j|$ and $\lambda_{0}:=\sum_{i,j=1}^{N_{d}}\lambda_{ij}|i\rangle\langle j|$. Note that since $\Sigma_{x}^{\dagger}=\Sigma_{x}$ and $\lambda_{0}^{\dagger}=\lambda_{0}$, for each $x$, we can find a unitary transformation $U_{x}$ that diagonalizes $\Sigma_{x}-\lambda_{0}$, giving us $\gamma_{i}^{x}\delta_{i\ell}=\sum_{j,k}[U^{\dagger}_{x}]_{ij}(\Sigma^{x}_{jk}-\lambda_{jk})[U_{x}]_{k\ell}$. Furthermore,
@@
where $F_{i}:=\sum_{j}K_{j}[U_{x}]_{ji}$ is at most $d$-local and $\sum_{i}F_{i}^{\dagger}F_{i}=I$. (Note that $F_{i}$ depends on $x$ in general, though here we suppress it to simplify the notation.) We therefore have $|\gamma_{i}^{x}|=\|F_{i}\|_{op}\cdot O(dN^{b-a})=O(dN^{b-a})$ from Lemma 2 and the fact that $\|F_{i}\|_{op}\leq 1$. This enables us to bound $\|\Sigma_{x}-\rho_{0}\|_{1}=\|\sum_{i=1}^{n_{d}}\gamma_{i}^{x}|i\rangle\langle i|\|_{1}=O(dn_{d}N^{b-a})$, which is independent of $x$. Using the Fuchs-van de Graaf inequality Eq. (S1), or $f(\Sigma_{x},\lambda_{0})\geq 1-\frac{1}{2}\|\Sigma_{x}-\lambda_{0}\|_{1}=1-O(n_{d}dN^{b-a})$, we have
$f((\Lambda+\mathcal{B})\otimes\mathrm{id}(|\psi\rangle\langle\psi|),\Lambda\otimes\mathrm{id}(|\psi\rangle\langle\psi|))$ $\geq\sum_{x=1}^{2^{k}}\langle x|B^{\dagger}B|x\rangle\cdot f\left(\sum_{i,j=1}^{n_{d}}\Sigma_{ij}^{x}|i\rangle\langle j|,\sum_{i,j=1}^{n_{d}}\lambda_{ij}|i\rangle\langle j|\right)$
@@
$\epsilon=D_{P}(\Lambda,\Lambda+\mathcal{B})=\sqrt{1-F(\Lambda,\Lambda+\mathcal{B})^{2}}=O(\sqrt{dn_{d}}N^{(b-a)/2})=O(\sqrt{d}(2s+1)^{d}N^{(b-a)/2})~{},$ (S19)
@@
Consider the code space $\mathfrak{C}=span\{|J\!,\!M\rangle;M\!=\!M_{\mathrm{min}},M_{\mathrm{min}}\!+\!\Delta,M_{\mathrm{min}}\!+\!2\Delta,\dots,M_{\mathrm{max}}\}$, where $J\sim N^{a}$, $M_{\mathrm{max}}\!-\!M_{\mathrm{min}}\sim N^{b}$, and $b<a\leq 1$. $\mathfrak{C}$ forms an $(\!(N,O(b\log_{2}N))\!)$ AQECC against a $d$-local noise channel with $\epsilon(d)=O(\sqrt{d}n_{d}N^{(b-a)/2})$ if we take $\Delta\geq 4sd+1$.

### v2

@@ Acknowledgements.
where $\lambda_{ij}$ are the components of a density operator, and $D_{P}(\Lambda+\mathcal{B},\Lambda)\leq\epsilon$, where $\Lambda(\rho)=\sum_{i,j}\lambda_{ij}\mathrm{Tr}(\rho)|i\rangle\langle j|$ and $(\Lambda+\mathcal{B})(\rho)=\Lambda(\rho)+\sum_{i,j}\mathrm{Tr}(\rho B_{ij})|i\rangle\langle j|$.
This theorem in fact has an interpretation from the complementary channel point of view, as described in Ref. [18]. More specifically, consider the noise channel $\mathcal{N}(\rho)=\sum_{i}K_{i}(\rho)K_{i}^{\dagger}$, whose complementary channel is $\hat{\mathcal{N}}(\rho)=\sum_{i,j}\mathrm{Tr}(K_{j}\rho K_{i}^{\dagger})|i\rangle\langle j|=(\Lambda+\mathcal{B})(\rho)$, where $|i\rangle\langle j|$ acts on the environment Hilbert space. If we have perfect Knill-Laflamme conditions, the complementary channel of the noise becomes $\hat{\mathcal{N}}(\rho)=\sum_{i,j}\lambda_{ij}|i\rangle\langle j|:=\lambda_{0}=\Lambda(\rho)$, which is a fixed density matrix $\lambda_{0}$ independent of the input $\rho$. In other words, the condition for having a perfect quantum error-correcting code is equivalent to the condition that the environment gains no information from the code space. And we indeed see that the AQECC generalization of the Knill-Laflamme conditions is a perturbation from the complementary channel point of view, where the environment can gain some $\epsilon$-small information from the code space. Note that, in the original theorem in Ref. [18], Bures distance $D_{B}(\mathcal{N},\mathcal{M})$ is used as the measure of channel distance, while here we opt for using the purified distance $D_{P}(\mathcal{N},\mathcal{M})$, where $D_{B}(\mathcal{N},\mathcal{M})^{2}\leq D_{P}(\mathcal{N},\mathcal{M})$ can be obtained using Eq. (S2).
Here we consider the noise channel $\mathcal{N}(\rho)=\sum_{i=1}^{n_{d}}K_{i}(\rho)K_{i}^{\dagger}$, where $n_{d}=(2s+1)^{2d}$ is the number of independent Kraus operators. We assume $K_{j}$ is at most $d$-local, and that all $K_{j}$ act on the same set of $d$ qudits, so $K_{i}^{\dagger}K_{j}$ is at most $d$-local as well. We are also interested in the regime $J\sim N^{a}$ and $(M_{\mathrm{max}}-M_{\mathrm{min}})\sim N^{b}$. From Lemma 1, for $|J,n\rangle,|J,m\rangle\in\mathfrak{C}$ and $n\neq m$, we have $\langle J\!,\!{n}|K_{i}^{\dagger}K_{j}|J\!,\!{m}\rangle=0$ if $\Delta\geq 2sd+1$. Now assuming we take $\lambda_{ij}=\langle J\!,\!M_{\mathrm{min}}|K^{\dagger}_{i}K_{j}|J\!,\!M_{\mathrm{min}}\rangle$,we then have $\Lambda(\rho)=\sum_{i,j=1}^{n_{d}}\lambda_{ij}\mathrm{Tr}[\rho]|i\rangle\langle j|$ and $(\Lambda+\mathcal{B})(\rho)=\sum_{i,j=1}^{n_{d}}\mathrm{Tr}[K_{j}\rho K_{i}^{\dagger}]|i\rangle\langle j|$, where the inputs of both channels are from the logical space.
@@
For any state $|\psi\rangle=\sum_{x,y}A_{xy}|x\rangle\otimes|y\rangle$, if we consider the operators $A=\sum_{x,y}A_{xy}|x\rangle\langle y|$ and $B:=A^{T}=\sum_{x,y}A_{xy}|y\rangle\langle x|$ and the state $|\Phi\rangle=\sum_{x=1}^{2^{k}}|x\rangle\otimes|x\rangle$, then we can express $|\psi\rangle=(A\otimes\mathrm{id})|\Phi\rangle=(\mathrm{id}\otimes B)|\Phi\rangle$. We can therefore rewrite $\mathcal{E}\otimes\mathrm{id}[(V\otimes I)(|\psi\rangle\langle\psi|)(V^{\dagger}\otimes I)]=(\mathrm{id}\otimes B)(\mathcal{E}\otimes\mathrm{id})(|\tilde{\Phi}\rangle\langle\tilde{\Phi}|)(\mathrm{id}\otimes B^{\dagger})$ for $\mathcal{E}=\Lambda$ and $\mathcal{E}=\Lambda+\mathcal{B}$, where we abbreviate $|\tilde{\Phi}\rangle:=(V\otimes I)|\Phi\rangle$. Moreover,
$(\Lambda+\mathcal{B})\otimes\mathrm{id}(|\tilde{\Phi}\rangle\langle\tilde{\Phi}|)=\sum_{x=1}^{2^{k}}\sum_{i,j=1}^{n_{d}}\langle J\!,\!M_{x}|K_{i}^{\dagger}K_{j}|J\!,\!M_{x}\rangle\cdot|i\rangle\langle j|\otimes|x\rangle\langle x|:=\sum_{x=1}^{2^{k}}\sum_{i,j=1}^{n_{d}}\Gamma^{x}_{ij}\cdot|i\rangle\langle j|\otimes|x\rangle\langle x|,$ (S15)
$\Lambda\otimes\mathrm{id}(|\tilde{\Phi}\rangle\langle\tilde{\Phi}|)=\sum_{x=1}^{2^{k}}\sum_{i,j=1}^{n_{d}}\langle J\!,\!M_{\mathrm{min}}|K_{i}^{\dagger}K_{j}|J\!,\!M_{\mathrm{min}}\rangle\cdot|i\rangle\langle j|\otimes|x\rangle\langle x|=\sum_{x=1}^{2^{k}}\sum_{i,j=1}^{n_{d}}\lambda_{ij}\cdot|i\rangle\langle j|\otimes|x\rangle\langle x|~,$ (S16)
where we have used Lemma 1 and denoted $\Gamma^{x}_{ij}=\langle J\!,\!M_{x}|K_{i}^{\dagger}K_{j}|J\!,\!M_{x}\rangle$.
From the normalization condition $\langle\psi|\psi\rangle=\sum_{x}\langle x|B^{\dagger}B|x\rangle=1$, we can interpret $\langle x|B^{\dagger}B|x\rangle$ as a probability. The fidelity can be rewritten as
@@
$=f\left(\sum_{x=1}^{2^{k}}\langle x|B^{\dagger}B|x\rangle\cdot\sum_{i,j=1}^{n_{d}}\Gamma_{ij}^{x}|i\rangle\langle j|\otimes\frac{B|x\rangle\langle x|B^{\dagger}}{\langle x|B^{\dagger}B|x\rangle},\sum_{x=1}^{2^{k}}\langle x|B^{\dagger}B|x\rangle\cdot\sum_{i,j=1}^{n_{d}}\lambda_{ij}|i\rangle\langle j|\otimes\frac{B|x\rangle\langle x|B^{\dagger}}{\langle x|B^{\dagger}B|x\rangle}\right)$
$\geq\sum_{x=1}^{2^{k}}\langle x|B^{\dagger}B|x\rangle\cdot f\left(\sum_{i,j=1}^{n_{d}}\Gamma_{ij}^{x}|i\rangle\langle j|\otimes\frac{B|x\rangle\langle x|B^{\dagger}}{\langle x|B^{\dagger}B|x\rangle},\sum_{i,j=1}^{n_{d}}\lambda_{ij}|i\rangle\langle j|\otimes\frac{B|x\rangle\langle x|B^{\dagger}}{\langle x|B^{\dagger}B|x\rangle}\right)$
$=\sum_{x=1}^{2^{k}}\langle x|B^{\dagger}B|x\rangle\cdot f\left(\sum_{i,j=1}^{n_{d}}\Gamma_{ij}^{x}|i\rangle\langle j|,\sum_{i,j=1}^{n_{d}}\lambda_{ij}|i\rangle\langle j|\right)~,$ (S17)
@@
We will use the Fuchs-van de Graaf inequality Eq. (S1) to bound the fidelity via the trace norm. To this end, we would like to bound $\|\Gamma_{x}-\lambda_{0}\|_{1}$, where $\Gamma_{x}:=\sum_{i,j=1}^{n_{d}}\Gamma_{ij}^{x}|i\rangle\langle j|$ and $\lambda_{0}:=\sum_{i,j=1}^{n_{d}}\lambda_{ij}|i\rangle\langle j|$. Note that since $\Gamma_{x}^{\dagger}=\Gamma_{x}$ and $\lambda_{0}^{\dagger}=\lambda_{0}$, for each $x$, we can find a unitary transformation $U_{x}$ that diagonalizes $\Gamma_{x}-\lambda_{0}$, giving us $\gamma_{i}^{x}\delta_{i\ell}=\sum_{j,k}[U^{\dagger}_{x}]_{ij}(\Gamma^{x}_{jk}-\lambda_{jk})[U_{x}]_{k\ell}$. Furthermore,
@@
where $F_{i}:=\sum_{j}K_{j}[U_{x}]_{ji}$ is at most $d$-local and $\sum_{i}F_{i}^{\dagger}F_{i}=I$. (Note that $F_{i}$ depends on $x$ in general, though here we suppress it to simplify the notation.) We therefore have $|\gamma_{i}^{x}|=\|F_{i}\|_{\text{op}}\cdot O(dN^{b-a})=O(dN^{b-a})$ from Lemma 2 and the fact that $\|F_{i}\|_{op}\leq 1$. This enables us to bound $\|\Gamma_{x}-\rho_{0}\|_{1}=\|\sum_{i=1}^{n_{d}}\gamma_{i}^{x}|i\rangle\langle i|\|_{1}=O(dn_{d}N^{b-a})$, which is independent of $x$. Using the Fuchs-van de Graaf inequality Eq. (S1), or $f(\Gamma_{x},\lambda_{0})\geq 1-\frac{1}{2}\|\Gamma_{x}-\lambda_{0}\|_{1}=1-O(n_{d}dN^{b-a})$, we have
$f((\Lambda+\mathcal{B})\otimes\mathrm{id}(|\psi\rangle\langle\psi|),\Lambda\otimes\mathrm{id}(|\psi\rangle\langle\psi|))$ $\geq\sum_{x=1}^{2^{k}}\langle x|B^{\dagger}B|x\rangle\cdot f\left(\sum_{i,j=1}^{n_{d}}\Gamma_{ij}^{x}|i\rangle\langle j|,\sum_{i,j=1}^{n_{d}}\lambda_{ij}|i\rangle\langle j|\right)$
@@
$\epsilon=D_{P}(\Lambda,\Lambda+\mathcal{B})=\sqrt{1-F(\Lambda,\Lambda+\mathcal{B})^{2}}=O(\sqrt{dn_{d}}N^{-(a-b)/2})=O(\sqrt{d}(2s+1)^{d}N^{-(a-b)/2})~,$ (S20)
@@
Consider the code space $\mathfrak{C}=span\{|J\!,\!M\rangle;M\!=\!M_{\mathrm{min}},M_{\mathrm{min}}\!+\!\Delta,M_{\mathrm{min}}\!+\!2\Delta,\dots,M_{\mathrm{max}}\}$, where $J\sim N^{a}$, $M_{\mathrm{max}}\!-\!M_{\mathrm{min}}\sim N^{b}$, and $b<a\leq 1$. $\mathfrak{C}$ forms an $(\!(N,O(\log_{2}\frac{N^{b}}{\Delta}))\!)$ AQECC against a $d$-local noise channel with $\epsilon(d)=O(\sqrt{d}n_{d}N^{(b-a)/2})$ if we take $\Delta\geq 4sd+1$.